Circular No. 8104

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2003 F2 (NEAT) An apparently asteroidal object reported by the NEAT project (discov-

ery observation given below), and posted on the NEO Confirmation Page, has been reported as faintly cometary by a few observers. G. Masi reports that CCD observations in good conditions (0′′.9 seeing) with the Danish 1.54-m telescope at the European Southern Observatory on Mar. 28.3 and 29.1 UT show the object to be nonstellar, with a slight elongation toward p.a. 315^{o}, such that a nuclear condensation appears on the southeast side of a coma that has size 5′′.5 along a southeast-northwest axis and 4′′ along a northeast-southwest axis. Images taken with the 1.06-m KLENOT telescope at Kleť on Mar. 31.9 by M. Tichý and M. Kočer show the object as slightly diffuse with a coma diameter of 6′′.

2003 UT  α_{2000}  δ_{2000}  m_{2}
Mar. 27.20121  12^{h}18^{m}44.^{s}52  −4^{o}22′58′′.2  20.4

The available astrometry, the following elliptical orbital elements, and an ephemeris appear on MPEC 2003-G04.

T = 2003 Apr. 24.365 TT  ω = 194.^{o}345
e = 0.54757  Ω = 358.782 }2000.0
q = 2.91198 AU  i = 11.107
a = 6.43630 AU n^{o} = 0.060360 P = 16.3 years

SUPERNOVAE 2003cu, 2003cv, AND 2003cw Further to IAUC 8103, W. M. Wood-Vasey, G. Aldering, and P. Nugent

report the discovery of three supernovae on unfiltered NEAT images:

SN  2003 UT  α_{2000}  δ_{2000}  Mag.  Offset
2003cu Mar. 27.14 10^{h}53^{m}02.^{s}66 − 0^{o}41′03′′.9 19.9 6′′.1 W, 3′′.6 N
2003cv Mar. 31.13 11 17 48.36 +19 09 08.5 17.6 0′′.6 W, 2′′.8 N
2003cw Mar. 31.13 11 24 03.88 +23 20 44.8 20.1 0′′.1 W, 10′′.0 N

Additional approximate CCD magnitudes: SN 2003cu, 2002 Apr. 5.14, [21.1; 2003 Feb. 23.54, 19.2; Mar. 9.36, 19.9; 24.16, 20.1; 26.14, 20.2; Apr. 1.22, 20.2. SN 2003cv, 2002 Apr. 1.33, [21.0; 2003 Mar. 24.18, 17.8; 26.29, 17.7. SN 2003cw, Mar. 8.46, [20.9; 24.18, 19.7.

2003 April 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green